OpenStudy (anonymous):

√3 sinx - 2sinx cosx = 0 sinx (√3 - 2cosx) = 0 Dont understand how this was factored?

OpenStudy (campbell_st):

ok so this ca be written as \[\sqrt{3} \times \sin(x) - 2 \times \sin(x) \cos(x) = 0\] what is common in both terms is sin(x) so it can be placed outside the brackets as he common factor \[\sin(x)(\sqrt{3} - 2\cos(x)) = 0\]

OpenStudy (campbell_st):

if you distribute in the normal way you will get the original equation....

OpenStudy (anonymous):

got it thanks